# Origen: Zero-Shot 3D Orientation Grounding in Text-to-Image Generation

### v1

@@ H Implementation Details
I More Qualitative Comparisons on Single Object Orientation Grounding
Orientation Zero-1-to-3 [46] C3DW [11] Origen (Ours) Orientation Zero-1-to-3 [46] C3DW [11] Origen (Ours)
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure C: Additional Qualitative results on the MS-COCO-Single benchmark.
Figure D: Qualitative comparisons on MS-COCO-4-View benchmark.
@@ I More Qualitative Comparisons on Single Object Orientation Grounding
In the following, Fig. C and Fig. D present additional qualitative comparisons on the MS-COCO-Single and MS-COCO-NView benchmarks.
@@
J More Qualitative Results on Multi Object Orientation Grounding
Orientation Origen (Ours) Orientation Origen (Ours) Orientation Origen (Ours) Orientation Origen (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure E: Additional Qualitative results on the MS-COCO-Multi benchmark.
@@
We report more qualitative results on multi object orientation grounding (MS-COCO-Multi) in Fig. E.

### v2

@@ Appendix I Computational Costs
Appendix J More Qualitative Comparisons on Single Object Orientation Grounding
Orientation Zero-1-to-3 [17] C3DW [18] FreeDoM [35] ReNO [22] Origen (Ours)
Figure 9: Qualitative comparisons on ORIBENCH-Single benchmark (Sec. 4.5). Compared to the existing orientation-to-image models [18, 17], Origen generates the most realistic images, which also best align with the grounding conditions in the leftmost column.
@@
In the following, Fig. 9 presents additional qualitative comparisons on the single object orientation grounding benchmarks.
@@
Appendix K More Qualitative Comparisons on Multi Object Orientation Grounding
Orientation DPS [32] MPGD [33] FreeDoM [35] ReNO [22] Origen (Ours)
Figure 10: Qualitative comparisons on ORIBENCH-Multi benchmark (Sec. 4.5). Compared to the guided-generation methods [32, 33, 35, 22], Origen generates the most realistic images, which also best align with the grounding conditions in the leftmost column.
@@
We report more qualitative results on multi object orientation grounding (ORIBENCH-Multi) in Fig. 10.